OpenStudy (anonymous):

In the following problem, verify that f(f^-1(x))=x and that f^-1 (f(x))= x

OpenStudy (anonymous):

\[f (x)= \left(\begin{matrix}6x-7 \\ 2-x\end{matrix}\right)\]

OpenStudy (anonymous):

and \[f^(-1) (x)= \left(\begin{matrix}2x+7 \\ 6+x\end{matrix}\right)\]

OpenStudy (anonymous):

f(x) wants a value x as input. Above, you've said you want that input to be f^-1(x). So, in f(x)=(6x-7)/(2-x) replace x with f^-1(x) = (2x+7)/(6+x). See what you get... shoudl come out as x. Now do similar by the other way...

OpenStudy (anonymous):

\[f(x)=\frac{6x-7}{2-x}\] \[f^{-1}(x)=\frac{2x+7}{6+x}\] \[f\circ f^{-1}(x)=f(f^{-1}(x))=\] \[=\frac{6\times (\frac{2x+7}{x+6})-7}{2-(\frac{2x+7}{6+x})}\]

OpenStudy (anonymous):

Thanks